# Optimal Pure Differentially Private Sparse Histograms in Near-Linear Deterministic Time

### v1

@@ Appendix C Proofs for Section 4
Case One: $J\nsubseteq S$, then
@@
Case Two: $J\subseteq S$ and $i^{*}\notin S$, then
$\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]$ $=0\leq\Pr\left[I=S\mid I_{1}=J\right].$ (23)
Case Three: $J\subseteq S$ and $i^{*}\in S$. First,
$\Pr\left[I=S\mid I_{1}=J\right]$ $=\Pr\left[I_{2}=S\setminus J\right]=\frac{1}{\binom{d-\lvert{J}\rvert}{n+k-\lvert{J}\rvert}},$ (24)
$\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]$ $=\Pr\left[I_{2}^{\prime}=S\setminus{\left({J\cup\{{i^{*}}\}}\right)}\right]=\frac{1}{\binom{d-\lvert{J}\rvert-1}{n+k-\lvert{J}\rvert-1}}.$ (25)
@@
$\frac{\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]missing}{\Pr\left[I=S\mid I_{1}=J\right]missing}$ $=\frac{\binom{a}{b}}{\binom{a-1}{b-1}}=\frac{a}{b}.$ (26)
@@
Proof of Lemma 4.4.
Let the $Z_{i}$ ($Z_{i}^{\prime}$) be defined as in Lemma 4.2. For each $J\subseteq\operatorname{\textsc{supp}}{({\vec{h}})}\cap\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}$, it holds that
$\Pr\left[I_{1}^{\prime}=J\right]$ $=\Pr\left[i^{*}\notin I_{1}^{\prime}\right]\cdot\prod_{i\in\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}\setminus\{{i^{*}}\}}\Pr\left[Z_{i}^{\prime}=\mathds{1}_{\left[i\in J\right]}\right]$
$=(1-p_{\tau})\cdot\prod_{i\in\operatorname{\textsc{supp}}{({\vec{h}})}\setminus\{{j^{*}}\}}\Pr\left[Z_{i}=\mathds{1}_{\left[i\in J\right]}\right]$
$=\Pr\left[j^{*}\notin I_{1}^{\prime}\right]\cdot\prod_{i\in\operatorname{\textsc{supp}}{({\vec{h}})}\setminus\{{j^{*}}\}}\Pr\left[Z_{i}=\mathds{1}_{\left[i\in J\right]}\right]=\Pr\left[I_{1}=J\right].$
Next,
$\Pr\left[I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]$ $=\Pr\left[i^{*}\in I_{1}^{\prime}\right]\cdot\prod_{i\in\operatorname{\textsc{supp}}{({\vec{h}^{\prime}})}\setminus\{{i^{*}}\}}\Pr\left[Z_{i}^{\prime}=\mathds{1}_{\left[i\in J\right]}\right]$
$=p_{\tau}\cdot\prod_{i\in\operatorname{\textsc{supp}}{({\vec{h}})}\setminus\{{j^{*}}\}}\Pr\left[Z_{i}=\mathds{1}_{\left[i\in J\right]}\right]$
$=\Pr\left[j^{*}\in I_{1}^{\prime}\right]\cdot\prod_{i\in\operatorname{\textsc{supp}}{({\vec{h}})}\setminus\{{j^{*}}\}}\Pr\left[Z_{i}=\mathds{1}_{\left[i\in J\right]}\right]=\Pr\left[I_{1}=J\cup\{{j^{*}}\}\right].$
Proof of Lemma 4.5.
$J\subseteq S$ and $i^{*}\notin S$, $j^{*}\in S$, then
$\Pr\left[I=S\mid I_{1}=J\right]$ $=\Pr\left[I_{2}=S\setminus J\right]=\frac{1}{\binom{d-\lvert{J}\rvert}{n+k-\lvert{J}\rvert}},$ (27)
$\Pr\left[I=S\mid I_{1}=J\cup\{{j^{*}}\}\right]$ $=\Pr\left[I_{2}=S\setminus{\left({J\cup\{{j^{*}}\}}\right)}\right]=\frac{1}{\binom{d-\lvert{J}\rvert-1}{n+k-\lvert{J}\rvert-1}},$ (28)
Denote $a=d-\lvert{J}\rvert$ and $b=n+k-\lvert{J}\rvert$.
$\frac{\Pr\left[I=S\mid I_{1}=J\cup\{{j^{*}}\}\right]missing}{\Pr\left[I=S\mid I_{1}=J\right]missing}$ $=\frac{\binom{a}{b}}{\binom{a-1}{b-1}}=\frac{a}{b}.$ (29)
@@ Appendix D Circuit-Based MPC Protocol for Sparse Histograms
In this section, we present a prototype secure multiparty computation (MPC) protocol for releasing an $\varepsilon$-DP sparse histogram with optimal $\ell_{\infty}$ error guarantees. Specifically, we provide a circuit-based implementation of Algorithm 3. Our design leverages the simple structure of Algorithm 3, which facilitates efficient circuit realization.
@@
    Add: Given $a,b\in\mathbb{N}$, outputs $a+b$. (Cost: $O(\omega)$)
@@
    Mul: Given $a,b\in\mathbb{N}$, outputs $a\cdot b$. (Cost: $O(\omega^{2})$)
@@
    LessThan: Given $a,b\in\mathbb{N}$, outputs $1$ if $a<b$, and 00 otherwise. (Cost: $O(\omega)$)
@@
    Equal: Given $a,b\in\mathbb{N}$, outputs $1$ if $a=b$, and 00 otherwise. (Cost: $O(\omega)$)
@@
    Mux: Given a bit $b\in\{0,1\}$ and values $x,y\in\mathbb{N}$, outputs $b\cdot x+(1-b)\cdot y$. (Cost: $O(\omega)$)
@@
    Sort: Given $a_{1},\ldots,a_{n}\in\mathbb{N}$, returns a sorted sequence $(b_{1},\ldots,b_{n})$. (Cost: $O(\omega\cdot n\ln n)$ (Ajtai et al., 1983))
@@
    RandomTableAccess: Given a public vector $\vec{a}$ and an index $J\in[m]$, returns $\vec{a}_{J}$. (Cost: $O(\omega\cdot m)$ for table of size $m$)
@@
    Clamp: Given $x\in\mathbb{N}$ and bounds $a,b\in\mathbb{N}$, outputs $\max(a,\min(x,b))$. This can be implemented using LessThan and Mux. (Cost: $O(\omega)$)
@@
    Bernoulli$(p)$: Outputs $1$ with probability $p$, and 00 otherwise. This can be implemented using a random input, LessThan, and Mux. (Cost: $O(\log(1/\delta))$ for precision $\delta$)
@@
D.1 Protocol
We will present a protocol achieving cost
$O{\left({\text{CostSort}{({n,\omega})}+n\cdot\omega+n\cdot\text{CostMul}{({\omega})}+n\cdot{\left({\frac{1}{\varepsilon}\cdot{\Big{(}{\ln\frac{1}{\varepsilon}+\ln\frac{d}{\gamma}}\Big{)}}+\ln^{2}{\Big{(}{\frac{d}{\varepsilon\gamma}}\Big{)}}}\right)}}\right)},$
where $\text{CostSort}{({n,\omega})}$ denotes the circuit cost of sorting $n$ elements, each represented with $\omega$ bits, and $\text{CostMul}{({\omega})}$ denotes the circuit cost of multiplying two $\omega$-bit numbers. The protocol is presented in Algorithm 8.
@@ D.1 Protocol
2: parameter $\tau\in\mathbb{N}_{+}$triangleright as defined in Equation 2
@@
9:         $c_{i}\leftarrow i-\textit{firstMinusOne}$triangleright compute frequency
@@
13:         Create dummy pair: (,)bottombottom($\bot$,$\bot$)( ⊥ , ⊥ )
@@
    If $Z<\tau$, convert the triple into a dummy triple (,,)bottombottombottom($\bot$,$\bot$,$\bot$)( ⊥ , ⊥ , ⊥ ).
@@
18:     elseCreate dummy triple: (,,)bottombottombottom($\bot$,$\bot$,$\bot$)( ⊥ , ⊥ , ⊥ )
@@
23:Convert each $i\in I_{2}$ to a triple $(i,0,2)$triangleright the label 2 indicates membership in $I_{2}$
@@
30:return the first $n+k$ pairs in $I$
@@
Given the participants’ data $x^{({1})},\ldots,x^{({n})}$, Algorithm 8 constructs an array $I_{1}$ consisting of all pairs ${({y,\vec{h}{\left[{y}\right]}})}$ for $y\in\operatorname{\textsc{supp}}(\vec{h})$, along with $n-\lvert{\operatorname{\textsc{supp}}(\vec{h})}\rvert$ dummy pairs to hide the true support size. This is achieved by first sorting the inputs $x^{({1})},\ldots,x^{({n})}$ into $y^{({1})},\ldots,y^{({n})}$. Then, for each $i\in[n]$, if $i$ corresponds to the last occurrence of $y^{({i})}$ in the sorted array, the algorithm creates a pair consisting of $y^{({i})}$ and the difference between $i$ and the index of $y^{({i})}$’s first appearance minus one—yielding its count $\vec{h}{\left[{y}\right]}$. Otherwise, the algorithm generates a dummy pair. Denote the array of the generated pairs as $I_{1}$.
@@
In this phase, for each $i\in[n]$, if $I_{1}[i]=(y,c)$ is a non-dummy pair, Algorithm 8 samples a noisy count $Z$ from the mechanism $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}(c)$ and creates a triple $(y,c,Z)$. If $Z<\tau$, the triple is replaced by a dummy triple (,,)bottombottombottom($\bot$,$\bot$,$\bot$)( ⊥ , ⊥ , ⊥ ). If $I_{1}[i]$ is a dummy pair, the algorithm directly creates a dummy triple. Finally, the algorithm collects all triples and sorts them so that non-dummy triples precede dummy ones. We reuse the notation and denote the resulting array as $I_{1}$ again.
After this step, the set of elements appearing in the first component of the triples in $I_{1}$ in Algorithm 8 exactly matches those in $I_{1}$ of Algorithm 3.
@@
Based on the proof of Theorem 5.2 and the structure of Algorithm 2, this mechanism can be implemented as follows:
  1. 1.
    Initialize an instance of $\mathcal{S}_{\operatorname{DLap}}{({\varepsilon,\delta})}$ as defined in Theorem 5.3, where $\delta=\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\cdot\frac{\varepsilon\gamma/d}{1-\varepsilon\gamma/d}\cdot\frac{1}{1+n}$.
  2. 2.
@@
    $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}(c)\doteq(1-B)\cdot\text{clamp}{[{{c+Z^{\prime}},{0},{n}}]}+B\cdot\operatorname{\mathbb{U}niform}\left({\mathbb{N}_{\leq{n}}}\right),$ (30)
@@
Under this setting, the Bernoulli variable $B\sim\operatorname{\mathbb{B}ernoulli}\left({\gamma^{\prime\prime}}\right)$ can be sampled exactly using the Bernoulli$(p)$ circuit primitive, which requires $O\left(\log_{2}\frac{d}{\varepsilon\gamma}\right)$ gates. Overall, the implementation of $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}(c)$ uses the following circuit components: one Bernoulli, one Mux, one Clamp, and one circuit for $\mathcal{S}_{\operatorname{DLap}}{({\varepsilon,\delta})}$.
@@
We briefly review the design of $\mathcal{S}_{\operatorname{DLap}}{({\varepsilon,\delta})}$ from Algorithm 6. Each sample $Z^{\prime}$ from this mechanism has the form
@@
    $X$ and $Y$ are sampled using the finite Alias method (Algorithm 5):
@@
As discussed in Section 4.2, to generate a uniformly random subset of size $n+k$ from $[d]$, we use the following efficient method:
@@
    Otherwise, terminate Algorithm 8 early and return a fixed sparse histogram, e.g., one in which elements $1$ through $n$ each have count $1$. While this may slightly degrade utility, it does not compromise privacy. As shown in Section 4.2, this increases the failure probability of the utility guarantee by at most an additive $\sqrt{n+k}\cdot\exp(-(n+k)/16)$.

### v2

@@ C.1. Privacy Guarantee
Case One
: $J\nsubseteq S$, then
@@
Case Two
: $J\subseteq S$ and $i^{*}\notin S$, then
(41) $\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]$ $=0\leq\Pr\left[I=S\mid I_{1}=J\right].$
Case Three
: $J\subseteq S$ and $i^{*}\in S$. First,
(42) $\Pr\left[I=S\mid I_{1}=J\right]$ $=\Pr\left[I_{2}=S\setminus J\right]=\frac{1}{\binom{d-\lvert{J}\rvert}{n+k-\lvert{J}\rvert}}$,
(43) $\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]$ $=\Pr\left[I_{2}^{\prime}=S\setminus{\left({J\cup\{{i^{*}}\}}\right)}\right]=\frac{1}{\binom{d-\lvert{J}\rvert-1}{n+k-\lvert{J}\rvert-1}}$.
@@
(44) $\frac{\Pr\left[I^{\prime}=S\mid I_{1}^{\prime}=J\cup\{{i^{*}}\}\right]}{\Pr\left[I=S\mid I_{1}=J\right]}$ $=\frac{\binom{a}{b}}{\binom{a-1}{b-1}}=\frac{a}{b}.$
@@ Appendix D Circuit-Based MPC Protocol for Sparse Histograms
In this section, we present a prototype secure multiparty computation (MPC) protocol for releasing an $\varepsilon$-DP sparse histogram with optimal $\ell_{\infty}$ error guarantees. Specifically, we provide a circuit-based implementation of Algorithm˜2. Our design leverages the simple structure of Algorithm˜2, which facilitates efficient circuit realization.
@@
    Add: Given $a,b\in\mathbb{N}$, outputs $a+b$. (Gates: $O(\omega)$)
@@
    Mul: Given $a,b\in\mathbb{N}$, outputs $a\cdot b$. (Gates: $O(\omega^{2})$)
@@
    LessThan: Given $a,b\in\mathbb{N}$, outputs $1$ if $a<b$, and 0 otherwise. (Gates: $O(\omega)$)
@@
    Equal: Given $a,b\in\mathbb{N}$, outputs $1$ if $a=b$, and 0 otherwise. (Gates: $O(\omega)$)
@@
    Mux: Given a bit $b\in\{0,1\}$ and values $x,y\in\mathbb{N}$, outputs $b\cdot x+(1-b)\cdot y$. (Gates: $O(\omega)$)
@@
    Sort: Given $a_{1},\ldots,a_{n}\in\mathbb{N}$, returns a sorted sequence $(b_{1},\ldots,b_{n})$. (Gates: $O(\omega\cdot n\ln n)$ (Ajtai et al., 1983))
@@
    RandomTableAccess: Given a public vector $\vec{a}$ and an index $J\in[m]$, returns $\vec{a}_{J}$. (Gates: $O(\omega\cdot m)$ for table of size $m$)
@@
    Clamp: Given $x\in\mathbb{N}$ and bounds $a,b\in\mathbb{N}$, outputs $\max(a,\min(x,b))$. This can be implemented using LessThan and Mux. (Gates: $O(\omega)$)
@@
    Bernoulli(p)(p): Outputs $1$ with probability $p$, and 0 otherwise. This can be implemented using a random input, LessThan, and Mux. (Gates: $O(\log(1/\delta))$ for precision $\delta$)
@@
Protocol
The protocol is presented in Algorithm˜7, with the main properties summarized below.
Theorem D.1 (Private Sparse Histogram Circuit).
Given integers $n,d,a_{\varepsilon},b_{\varepsilon},a_{\gamma},b_{\gamma}\in\mathbb{N}_{+}$ that fit in a constant number of machine words, and $\gamma\doteq a_{\gamma}/b_{\gamma}\in{({1/n^{O(1)},1})}$, there exists a circuit with cost
$O{\left({\text{CostSort}{({n,\omega})}+n\cdot\omega+n\cdot\text{CostMul}{({\omega})}+n\cdot{\Big({\frac{1}{\varepsilon}\cdot{\Big({\ln\frac{1}{\varepsilon}+\ln\frac{d}{\gamma}}\Big)}+\ln^{2}{\Big({\frac{d}{\varepsilon\gamma}}\Big)}}\Big)}}\right)},$
where $\text{CostSort}{({n,\omega})}$ is the circuit cost of sorting $n$ elements represented using $\omega$ bits, and $\text{CostMul}{({\omega})}$ is the circuit cost of multiplying two $\omega$-bit numbers. This circuit implements a mechanism that takes a dataset $\mathcal{X}$ of $n$ user-contributed elements from domain $[d]$ as input, and outputs a histogram $\tilde{h}\in\mathbb{N}_{\leq{n}}^{d}$ approximating the original histogram $\vec{h}$ of $\mathcal{X}$, such that $\lVert{\tilde{h}}\rVert_{0}\in O(n)$ and
  • $\triangleright$
    Privacy Guarantee: $\mathcal{M}_{\textsc{\scriptsize Hist}}$ satisfies $2\varepsilon$-differential privacy, where $\varepsilon\doteq a_{\varepsilon}/b_{\varepsilon}\in\mathbb{Q}_{+}$.
  • $\triangleright$
    Utility Guarantee: There exists a universal constant $c_{\alpha}\in\mathbb{R}_{+}$ such that for every $\beta\geq 2\varepsilon\gamma$, letting $\alpha\doteq(c_{\alpha}/\varepsilon)\cdot\ln{({d/\beta})}$, the output $\tilde{h}$ is an $(\alpha,\beta)$-simultaneous accurate estimator of $\vec{h}$.
@@
1:Input: parameters $d,n,k=3n,a_{\varepsilon},b_{\varepsilon},a_{\gamma},b_{\gamma}\in\mathbb{N}_{+}$, s.t., $\varepsilon=a_{\varepsilon}/b_{\varepsilon}$ and $\gamma=a_{\gamma}/b_{\gamma}$
2: parameter $\tau\in\mathbb{N}_{+}$ $\triangleright$ as defined in Equation˜2
@@
9:   $c_{i}\leftarrow i-\textit{firstMinusOne}$ $\triangleright$ compute frequency
@@
13:   Create dummy pair: $(\bot,\bot)$
@@
    If $Z<\tau$, convert the triple into a dummy triple $(\bot,\bot,\bot)$.
@@
18:  elseCreate dummy triple: $(\bot,\bot,\bot)$
@@
23:Convert each $i\in I_{2}$ to a triple $(i,0,2)$ $\triangleright$ the label 2 indicates membership in $I_{2}$
@@
30:return the first $n+k$ pairs in II
@@
Given the participants’ data $x^{({1})},\ldots,x^{({n})}$, Algorithm˜7 constructs an array $I_{1}$ consisting of all pairs $({y,\vec{h}{\left[{y}\right]}})$} for $y\in\operatorname{\textsc{supp}}(\vec{h})$, along with $n-\lvert{\operatorname{\textsc{supp}}(\vec{h})}\rvert$ dummy pairs to hide the true support size. This is achieved by first sorting the inputs $x^{({1})},\ldots,x^{({n})}$ into $y^{({1})},\ldots,y^{({n})}$. Then, for each $i\in[n]$, if $i$ corresponds to the last occurrence of $y^{({i})}$ in the sorted array, the algorithm creates a pair consisting of $y^{({i})}$ and the difference between $i$ and the index of $y^{({i})}$’s first appearance minus one—yielding its count $\vec{h}{\left[{y}\right]}$. Otherwise, the algorithm generates a dummy pair. Denote the array of the generated pairs as $I_{1}$.
@@
In this phase, for each $i\in[n]$, if $I_{1}[i]=(y,c)$ is a non-dummy pair, Algorithm˜7 samples a noisy count $Z$ from the mechanism $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}(c)$ and creates a triple $(y,c,Z)$. If $Z<\tau$, the triple is replaced by a dummy triple $(\bot,\bot,\bot)$. If $I_{1}[i]$ is a dummy pair, the algorithm directly creates a dummy triple. Finally, the algorithm collects all triples and sorts them so that non-dummy triples precede dummy ones. We reuse the notation and denote the resulting array as $I_{1}$ again.
After this step, the set of elements appearing in the first component of the triples in $I_{1}$ in Algorithm˜7 exactly matches those in $I_{1}$ of Algorithm˜2.
@@
Based on the proof of Theorem˜B.3 and the structure of Algorithm˜1, this mechanism can be implemented as follows:
  1. (1)
    Initialize an instance of $\mathcal{S}_{\operatorname{DLap}}{({\varepsilon,\delta})}$ as defined in Theorem˜B.4, where $\delta=\frac{e^{\varepsilon}-1}{e^{\varepsilon}+1}\cdot\frac{\varepsilon\gamma/d}{1-\varepsilon\gamma/d}\cdot\frac{1}{1+n}$.
  2. (2)
@@
    (45) $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}(c)\doteq(1-B)\cdot\text{clamp}{[{{c+Z^{\prime}},{0},{n}}]}+B\cdot\operatorname{\mathbb{U}niform}\left({\mathbb{N}_{\leq{n}}}\right),$
@@
Under this setting, the Bernoulli variable $B\sim\operatorname{\mathbb{B}ernoulli}\left({\gamma^{\prime\prime}}\right)$ can be sampled exactly using the Bernoulli(p)(p) circuit primitive, which requires $O\left(\log_{2}\frac{d}{\varepsilon\gamma}\right)$ gates. Overall, the implementation of $\mathcal{M}_{\operatorname{\mathbb{PA}px\mathbb{DL}ap}{({n,\varepsilon,\varepsilon\gamma/d})}}(c)$ uses the following circuit components: one Bernoulli, one Mux, one Clamp, and one circuit for $\mathcal{S}_{\operatorname{DLap}}{({\varepsilon,\delta})}$.
@@
We briefly review the design of $\mathcal{S}_{\operatorname{DLap}}{({\varepsilon,\delta})}$ from Algorithm˜6. Each sample $Z^{\prime}$ from this mechanism has the form
@@
    $X$ and $Y$ are sampled using the finite Alias method (Algorithm˜5):
@@
As discussed in Section˜4.2, to generate a uniformly random subset of size $n+k$ from $[d]$, we use the following efficient method:
@@
    Otherwise, terminate Algorithm˜7 early and return a fixed sparse histogram, e.g., one in which elements $1$ through $n$ each have count $1$. While this may slightly degrade utility, it does not compromise privacy. As shown in Section˜4.2, this increases the failure probability of the utility guarantee by at most an additive $\sqrt{n+k}\cdot\exp(-(n+k)/16)$.